Quick Start Company makes 12-volt car batteries. After many years of product testing, the company knows that the average life of a Quick Start battery is normally distributed, with a mean of 45.8 months and a standard deviation of 6.9 months. (a) If Quick Start guarantees a full refund on any battery that fails within the 36-month period after purchase, what percentage of its batteries will the company expect to replace? (Round your answer to two decimal places.) %

(b) If Quick Start does not want to make refunds for more than 14% of its batteries under the full-refund guarantee policy, for how long should the company guarantee the batteries (to the nearest month)? months
Answer #1

a) P(X < 36)

= P(z < (36 - 45.8)/6.9)

= P(z < -1.42)

= 0.0778

= 7.78 %

b) z score for bottom 14% area = -1.08

Hence,

Guarantee time period = 45.8 - 1.08*6.9 = 38